OpenStudy (anonymous):

How many license plates can be made which have the form ##LLL (2 numbers followed by 3 letters with repetition allowed) ? A. 1,757,600 B. 175,600 C. 3,840,500 D. 2,947,000

OpenStudy (anonymous):

10C2 times 26C3

OpenStudy (anonymous):

ohhh so it would be A?

OpenStudy (anonymous):

I dunno. use a calculator to find out.

OpenStudy (anonymous):

lol i did :)

OpenStudy (anonymous):

Thanks :) helpful very helpful!

OpenStudy (anonymous):

heh

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

I'm not sure that's right because 10C2 doesn't allow for repetition.

OpenStudy (anonymous):

hold on. is it with repetition?

OpenStudy (anonymous):

(2 numbers followed by 3 letters with repetition allowed) ?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

So then that WOULD be the answer correct?

OpenStudy (anonymous):

right, my bad. so it 2 numbers followed by letters with repetition. that means the letters can repeat, but not the numbers.

OpenStudy (anonymous):

... I think the numbers and letters can repeat

OpenStudy (anonymous):

I.E. you can have 00 or 99, etc. and you can have LLL or AAA, etc.

OpenStudy (anonymous):

\[10\times 10 \times 26\times 26 \times26\]

OpenStudy (anonymous):

yeah I think it's A

OpenStudy (anonymous):

yeah it is A

OpenStudy (anonymous):

that is, if repetition is allowed for both numbers and letters, and it does look like it is allowed for both numbers and letters.

OpenStudy (anonymous):

thought so :) Yes im pretty sure it is :)

OpenStudy (gw2011):

The correct answer is A. Since repetition is allowed, then you have 10 numbers and 26 letters to choose in order to place in the positions indicated. Therefore, the solution is as follows: (10^2)(26^3) = 1,757,600

OpenStudy (anonymous):

Ahhh alrighty :) Thanks

OpenStudy (gw2011):

You're welcome